# A Survey on 3D Gaussian Splatting

### v3

@@ 2 Background
In this section, we first provide a brief formulation of radiance fields (Sec. 2.1), a key concept in scene rendering. It outlines two main types of radiance field representations: implicit, like NeRF [12], which uses neural networks for a straightforward yet computationally demanding rendering; and explicit, like grid [25], which employs discrete structures for quicker access but at the cost of higher memory use. Sec. 2.2 further establishes linkages with relevant fields such as scene reconstruction and rendering. For a comprehensive overview of radiance fields, scene reconstruction and representation, and rendering methods, please see the excellent surveys [26, 27, 28, 29, 30, 31] for more insights.
@@ 3.2.2 Density Control
Though 3D GS has achieved impressive milestones, significant room for improvement remains, e.g., data and hardware requirement, rendering and optimization algorithm, and applications in downstream tasks. In the subsequent sections, we seek to elaborate on select extended versions of 3D GS. These are: i) Data-efficient 3D GS [46, 47, 48, 49, 50, 51, 52, 53, 54, 55] (Sec. 4.1), ii) Memory-efficient 3D GS [56, 57, 58, 59, 60, 61, 62, 63, 64] (Sec. 4.2), iii) Photorealistic 3D GS [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80] (Sec. 4.3), iv) Improved Optimization Algorithms [81, 22, 82, 77, 83, 84, 85, 86] (Sec. 4.4), v) 3D Gaussian with More Properties [87, 88, 89, 90, 91, 92, 93] (Sec. 4.5), and vi) 3D GS with Structured Information [94, 95, 96] (Sec. 4.6).
@@ 4.1 Data-efficient 3D GS
Two primary strategies are used for data-efficient 3D GS. i) Regularization based methods introduce additional constraints such as depth information to enhance the detail and global consistency [47, 50, 52]. For example, DNGaussian [50] introduced a depth-regularized approach to address the challenge of geometry degradation in sparse input views through hard and soft depth regularization and global-local depth normalization. It rectified the geometry using depth information extracted from pre-trained monocular depth estimators and optimized the positioning of Gaussian primitives without changing their shape. Unfortunately, when dealing with a limited number of views, or even just one, the efficacy of regularization techniques tends to diminish, which leads to ii) generalizability based methods that focus in particular on learning priors [48, 49, 54]. One typical implementation is to generate 3D Gaussians that can be used for rendering directly without optimization, by using deep neural networks. This paradigm typically requires multiple views for training but can reconstruct the 3D scene with only one input image. For instance, PixelSplat [48] proposed to sample Gaussians from dense probability distributions. It incorporated a multi-view epipolar transformer and a reparameterization trick to avoid local minima and maintain gradient flow. Splatter Image [49] applied GS in a monocular setting through a learning-based approach, utilizing a 2D image-to-image network that maps an input image to a 3D Gaussian per pixel. Note that this paradigm is mainly focused on the reconstruction of objects, and its generalizability leaves a lot of room for improvement.
@@ 5.3 AI-Generated Content (AIGC)
This explicit nature of 3D GS facilitates real-time rendering capabilities and unprecedented levels of control and editability, making it highly relevant for and AIGC applications. The explicit scene representations and differentiable rendering algorithms of 3D GS align perfectly with the requirements of AIGC for generating high-fidelity, real-time, and editable content, crucial for applications in virtual reality, interactive media, and beyond. Recent works have effectively utilized 3D GS in conjunction with fields like generative models [182, 126, 127, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 128, 193, 194, 195, 129, 196, 197, 198, 199, 200, 201, 202, 203, 204], avatars [205, 206, 207, 130, 208, 209, 210, 211, 212, 213, 214, 131, 215, 132, 216, 217, 23, 218, 219, 220, 221, 222, 223, 224, 225, 226, 133, 227], and scene editing [134, 135, 228, 91, 92, 90, 229, 121, 120, 119, 230, 231, 232, 233, 234, 235, 236, 136, 237, 102]. For instance, DreamGaussian [126] accelerated the generation of photorealistic 3D assets from single-view image through a three-step process: a diffusion based generative GS process, followed by an efficient algorithm for mesh extraction from 3D Gaussians based on local density querying, and finally a UV-space refinement stage to improve texture details. By combining 3D GS with a parametric morphable face model, GaussianAvatars [213] offered enhanced fidelity and flexibility in avatar animation, significantly improving upon current methods in novel-view rendering and expression reenactment. To improve the efficacy of text instruction based editing, Chen et al. [134] devised a semantic tracker to track the editing target during training while Fang et al. [135] proposed to extract the region of interest corresponding to the instruction. These advancements hold promise for numerous industrial applications, including digital assetization (with mesh extraction), the generation of long-form videos (e.g., Sora), etc.
@@ 5.6 Large-scale Scene Reconstruction
To address the highlighted challenges, researchers have made significant strides in two key areas: i) For training (learning 3D Gaussians for a scene), a divide-and-conquer strategy [146, 147, 148, 145] has been adopted, which segments a large scene into multiple, independent cells. This facilitates parallel optimization for expansive environments. An additional challenge lies in maintaining visual quality, as large-scale scenes often feature texture-less surfaces that can hamper the effectiveness of optimization such as Gaussian initialization and density control (Sec. 3.2). Enhancing the optimization algorithm presents a viable solution to mitigate this issue [45, 147]. ii) Regarding rendering, the adoption of the Level of Details (LoD) technique from computer graphics has proven instrumental. This technique adjusts the complexity of 3D scenes to balance visual quality with computational efficiency. Current implementations involve feeding only the essential Gaussians to the rasterizer [147], or designing explicit LoD structures like the Octree [148] and hierarchy [148]. Furthermore, integrating extra input modalities like LiDAR can further enhanced the reconstruction process [149, 150, 151]. One prominent challenge in large-scale scene reconstruction is limited capture, which could be alleviated by utilizing good priors (see “learning physical priors from large-scale data.” in Sec. 7).
@@ 7 Future Research Directions
$\bullet$ Empowering 3D GS with More Possibilities. Despite the significant potential of 3D GS, the full scope of applications for 3D GS remains largely untapped. A promising avenue for exploration involves augmenting 3D Gaussians with additional attributes (e.g., linguistic and spatiotemporal properties as mentioned in Sec. 4.5) and introducing structured information (e.g., spatial MLPs and grids as mentioned in Sec. 4.6), tailored for specific applications. Moreover, recent studies have begun to unveil the capability of 3D GS in several domains, e.g., the capture of hand-object interactions [285], the quantification of uncertainty [286], and fluid synthesis [287]. These preliminary findings highlight a significant opportunity for interdisciplinary scholars to explore 3D GS further.

### v4

@@ 2 Background
In this section, we first provide a brief formulation of radiance fields (Sec. 2.1), a key concept in scene rendering. It outlines two main types of radiance field representations: implicit, like NeRF [12], which uses neural networks for a straightforward yet computationally demanding rendering; and explicit, like grid [25], which employs discrete structures for quicker access but at the cost of higher memory use. Sec. 2.2 further establishes linkages with relevant fields such as scene reconstruction and rendering. For a comprehensive overview of radiance fields, scene reconstruction and representation, and rendering methods, please see the excellent surveys [26, 27, 28, 29, 30] for more insights.
@@ 3.2.2 Density Control
Though 3D GS has achieved impressive milestones, significant room for improvement remains, e.g., data and hardware requirement, rendering and optimization algorithm, and applications in downstream tasks. In the subsequent sections, we seek to elaborate on select extended versions of 3D GS. These are: i) Data-efficient 3D GS [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] (Sec. 4.1), ii) Memory-efficient 3D GS [56, 57, 58, 59, 60, 61, 62, 63, 64] (Sec. 4.2), iii) Photorealistic 3D GS [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80] (Sec. 4.3), iv) Improved Optimization Algorithms [81, 22, 82, 77, 83, 84, 85, 86] (Sec. 4.4), v) 3D Gaussian with More Properties [87, 88, 89, 90, 91, 92, 93] (Sec. 4.5), and vi) 3D GS with Structured Information [94, 95, 96] (Sec. 4.6).
@@ 4.1 Data-efficient 3D GS
Two primary strategies are used for data-efficient 3D GS. i) Regularization based methods introduce additional constraints such as depth information to enhance the detail and global consistency [49, 46, 51, 55]. For example, DNGaussian [49] introduced a depth-regularized approach to address the challenge of geometry degradation in sparse input views. FSGS [46] devised a Gaussian Unpooling process for initialization and also introduced depth regularization. MVSplat [51] proposed a cost volume representation so as to provide geometry cues. Unfortunately, when dealing with a limited number of views, or even just one, the efficacy of regularization techniques tends to diminish, which leads to ii) generalizability based methods that focus specially on learning priors [47, 48, 53]. One typical implementation is to generate 3D Gaussians that can be used for rendering directly without optimization, by using deep neural networks. This paradigm typically requires multiple views for training but can reconstruct the 3D scene with only one input image. For instance, PixelSplat [47] proposed to sample Gaussians from dense probability distributions. It incorporated a multi-view epipolar transformer and a reparameterization trick to avoid local minima and maintain gradient flow. Splatter Image [48] applied GS in a monocular setting through a learning-based approach, utilizing a 2D image-to-image network that maps an input image to a 3D Gaussian per pixel. Note that this paradigm is mainly focused on the reconstruction of objects, and its generalizability leaves a lot of room for improvement.
@@ 5.3 AI-Generated Content (AIGC)
This explicit nature of 3D GS facilitates real-time rendering capabilities and unprecedented levels of control and editability, making it highly relevant for and AIGC applications. The explicit scene representations and differentiable rendering algorithms of 3D GS align perfectly with the requirements of AIGC for generating high-fidelity, real-time, and editable content, crucial for applications in virtual reality, interactive media, and beyond. Recent works have effectively utilized 3D GS in conjunction with fields like generative models [183, 126, 127, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 128, 194, 195, 196, 129, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207], avatars [208, 209, 130, 210, 211, 212, 213, 214, 215, 131, 216, 132, 217, 218, 23, 219, 220, 221, 222, 223, 224, 225, 226, 227, 133, 228, 229], and scene editing [134, 135, 230, 91, 92, 90, 231, 121, 120, 119, 232, 233, 234, 235, 236, 237, 238, 136, 239, 102]. For instance, DreamGaussian [126] accelerated the generation of photorealistic 3D assets from single-view image through a three-step process: a diffusion based generative GS process, followed by an efficient algorithm for mesh extraction from 3D Gaussians based on local density querying, and finally a UV-space refinement stage to improve texture details. By combining 3D GS with a parametric morphable face model, GaussianAvatars [214] offered enhanced fidelity and flexibility in avatar animation, significantly improving upon current methods in novel-view rendering and expression reenactment. To improve the efficacy of text instruction based editing, Chen et al. [134] devised a semantic tracker to track the editing target during training while Fang et al. [135] proposed to extract the region of interest corresponding to the instruction. These advancements hold promise for numerous industrial applications, including digital assetization (with mesh extraction), the generation of long-form videos (e.g., Sora), etc.
@@ 5.6 Large-scale Scene Reconstruction
To address the highlighted challenges, researchers have made significant strides in two key areas: i) For training, a divide-and-conquer strategy [146, 147, 148, 145] has been adopted, which segments a large scene into multiple, independent cells. This facilitates parallel optimization for expansive environments. With the same spirit, Zhao et al. [152] proposed a distributed implementation of 3D GS training. An additional challenge lies in maintaining visual quality, as large-scale scenes often feature texture-less surfaces that can hamper the effectiveness of optimization such as Gaussian initialization and density control (Sec. 3.2). Enhancing the optimization algorithm presents a viable solution to mitigate this issue [44, 147]. ii) Regarding rendering, the adoption of the Level of Details (LoD) technique from computer graphics has proven instrumental. LoD adjusts the complexity of 3D scenes to balance visual quality with computational efficiency. Current implementations involve feeding only the essential Gaussians to the rasterizer [147], or designing explicit LoD structures like the Octree [148] and hierarchy [145]. Furthermore, integrating extra input modalities like LiDAR can further enhanced the reconstruction process [149, 150, 151]. One prominent challenge in large-scale scene reconstruction is limited capture, which could be alleviated by utilizing good priors (see “learning physical priors from large-scale data.” in Sec. 7).
@@ 7 Future Research Directions
$\bullet$ Empowering 3D GS with More Possibilities. Despite the significant potential of 3D GS, the full scope of applications for 3D GS remains largely untapped. A promising avenue for exploration involves augmenting 3D Gaussians with additional attributes (e.g., linguistic and spatiotemporal properties as mentioned in Sec. 4.5) and introducing structured information (e.g., spatial MLPs and grids as mentioned in Sec. 4.6), tailored for specific applications. Moreover, recent studies have begun to unveil the capability of 3D GS in several domains, e.g., point cloud registration [287], image representation and compression [60], and fluid synthesis [288]. These findings highlight a significant opportunity for interdisciplinary scholars to explore 3D GS further.